Find the inverse of the function.
f(x)=2x+1

Answer by nerdybill(6958)

f(x)=2x+1
y = 2x+1
switch the x and y's:
x = 2y+1
solve for y:
x = 2y+1
x-1 = 2y
x/2 - 1/2 = y
.
therefore
f'(x) = x/2 - 1/2
or
f'(x) = (1/2)x - 1/2